# Exploring Self-Interacting Dark Matter Halos with Diverse Baryonic Distributions: A Parametric Approach

### v1

@@ I Introduction
Cold dark matter (CDM) aligns well with observations at scales larger than megaparsecs. However, discrepancies arise at smaller scales in the inner structures of dark matter (DM) halos, where the density profiles can be either cored or cuspy Flores and Primack (1994); Spergel and Steinhardt (2000); de Blok et al. (2001); Kuzio de Naray et al. (2010); Oh et al. (2011); Oman et al. (2015). The cored cases may exhibit fainter or more diffuse profiles than most models predicted, while the cuspy cases can display steeper inner-density slopes. To explain these variations, researchers have elaborated on both baryonic feedback models and self-interacting dark matter (SIDM) models. Baryonic effects diversify the inner halo structures because it alters the structure by leading to compact stellar distributions through dissipation while simultaneously dispersing these distributions through feedback. The effectiveness of these processes varies with scale and depends on complex mechanisms including star formation and supernova feedback Navarro et al. (1996); Gnedin and Zhao (2002); Read and Gilmore (2005); Governato et al. (2010); Di Cintio et al. (2014); Chan et al. (2015); Tollet et al. (2016); Lazar et al. (2020); Macciò et al. (2020); Santos-Santos et al. (2020a). Similarly, SIDM contributes to diversity in inner halo structures through elastic scatterings that facilitate energy transfer within halos. This results in core formation and core collapse stages in the thermodynamic evolution of self-gravitating halos, known as gravothermal evolution Tulin and Yu (2018); Adhikari et al. (2022); Spergel and Steinhardt (2000); Kochanek and White (2000); Kamada et al. (2017); Ren et al. (2019); Santos-Santos et al. (2020b); Spergel and Steinhardt (2000); Correa et al. (2022); Yang et al. (2023a); Balberg et al. (2002, 2002); Koda and Shapiro (2011); Nishikawa et al. (2020); Sameie et al. (2020); Zavala et al. (2019); Kaplinghat et al. (2019); Kahlhoefer et al. (2019); Correa (2021); Turner et al. (2021); Slone et al. (2023); Silverman et al. (2022); Correa et al. (2022); Yang et al. (2023a); Minor et al. (2021); Yang and Yu (2021); Gilman et al. (2021, 2023); Nadler et al. (2023); Pollack et al. (2015); Choquette et al. (2019); Feng et al. (2021, 2022); Meshveliani et al. (2023); Fischer et al. (2024); Dhanasingham et al. (2023); Zhang et al. (2024). In probing SIDM, the diversity from baryon physics naturally integrate through gravitational interactions, resulting in a nonlinear and complex combined evolution. Nonetheless, understanding the interplay between baryons and dark matter is vital for making realistic SIDM predictions, as we depend on visible baryonic matter to probe the elusive dark matter structures Vogelsberger et al. (2014); Robertson et al. (2019, 2021); Yang and Yu (2021); Rose et al. (2023); Mastromarino et al. (2023); Rahimi et al. (2023); Zhong et al. (2023); Creasey et al. (2017); Fry et al. (2015); Kong et al. (2024); Fischer et al. (2023).
@@
To date, there have been only a handful of hydrodynamical cosmological simulations that incorporate SIDM physics Robertson et al. (2018); Robles et al. (2017); Elbert et al. (2018); Despali et al. (2019); Robertson et al. (2019). These simulations have revealed a complex interplay dependent on mass, galaxy type, and accretion history. Meanwhile, theoretical studies have used conducting fluid models, semi-analytic models, and idealized N-body simulations to model the combined effects of a baryonic potential and gravothermal evolution Balberg et al. (2002); Koda and Shapiro (2011); Essig et al. (2019); Nishikawa et al. (2020); Feng et al. (2021); Yang et al. (2023b); Jiang et al. (2023); Yang et al. (2023c); Zhong et al. (2023). It has been illuminated that galaxy formation can adiabatically contract dark matter, increase the DM scattering rate, and thus accelerate the gravothermal evolution. Some empirical relations have also been proposed to enhance our understanding of these complex processes Elbert et al. (2018); Sameie et al. (2018); Feng et al. (2021); Yang et al. (2023c); Zhong et al. (2023).
@@
In Fig. 1, we highlight the diverse baryonic structures across a wide range of samples depicted on the plane of scale density ratio $\rho_{H}/\rho_{s}$ versus scale radius ratio $r_{H}/r_{s}$. Here, the halos and their stellar components are modeled using NFW and Hernquist Hernquist (1990) profiles, characterized by scale parameters $\rho_{s}$, $r_{s}$ for halos and $\rho_{H}$, $r_{H}$ for stellar components. The color codes on the plot correspond to the stellar-to-halo mass ratio, represented on a logarithmic scale, revealing a trend where higher $M_{s}/M_{h}$ halos move up along the $\ln(y)\propto\ln(y)$ direction, exhibiting larger $\rho_{H}/\rho_{s}$ and $r_{H}/r_{s}$ ratios. Red dots indicate benchmark halos from Table 1, utilized for validating the model. These halos cover a broad spectrum in the $\rho_{H}/\rho_{s}$ and $r_{H}/r_{s}$ plane, demonstrating the model’s broad applicability.
@@
Table 1: Properties of N-body simulated benchmarks: The first four simulations, introduced in Ref.Zhong et al. (2023), correspond to the initial block of cases in Table 1 therein. The final two benchmarks, DM13+baryon2 and DM13 extreme are analogous to J1636+4707 and JWST-ER1, as described in Ref. Kong et al. (2024). Additionally, the DM12+baryon, DM13 compact, and DM13+baryon benchmarks represent further simulations conducted to broaden the range of benchmark scenarios. The core collapse times are computed considering $\sigma/m=10~{}\rm cm^{2}/g$ for the DM11 cases in the first block and $\sigma/m=1~{}\rm cm^{2}/g$ for the DM12 and DM13 cases, to be consistent with the curves in Fig. 4 (left panel). We use $t_{c,b}$ and $t_{c,0}$ to represent the core collapse times in the DM+baryon and DM-only scenarios. The halo masses are calculated as the mass enclosed within radii where the density is 200 times the critical density ($R_{\rm 200}$), using the scale density ($\rho_{s}$) and scale radius ($r_{s}$) of the NFW profile. The stellar masses are total masses computed using the scale density ($\rho_{H}$) and scale radius ($r_{H}$) of the Hernquist profile.
@@ III.2 A universal form factor for normalizing the evolution time with baryons
Figure 2: The unbroken universality in the evolution timeline with baryons, depicted through the evolution of normalized density profiles. For clarity, three benchmark cases—DM11 (blue), DM11+baryon (red), and DM13 compact (gray)—from Table 1 are illustrated to demonstrate the predictions of the parametric model. All the cases evolve into a fixed configuration as $t/t_{c}$ approaching $1.1$. For more details, refer to Section V and Fig. 4.
Before delving into the specifics, we illustrate in Fig. 2 the aspects just discussed. In this figure, the DM-only case is depicted in blue. The other cases include different amounts of baryons, and their inner density evolution and density profiles are compared with the DM-only case in the left and right panels, respectively. Both panels are normalized to depict the universality and its breaking clearly. From this figure, it is evident that all scenarios follow a unified evolution timeline, represented as $t/t_{c}$ and ranging from 00 to $1.1$. We define each point along this timeline as a “gravothermal phase.” The density profiles, conversely, vary according to the amount of baryonic matter and show different deviations from the DM-only scenario. Subsequently, we will discuss methods to normalize the evolution time and adjust the DM density profile for a parametric model.
@@
Figure 3: Contours of ${\cal F}_{t}\equiv t_{c,b}/t_{c,0}$ on the $\rho_{H}/\rho_{s}$ ($\hat{\rho}_{H}$) versus $r_{H}/r_{s}$ ($\hat{r}_{H}$) plane, depicted together with the isolated halos in the TNG-50-1 simulation (black dots) Nelson et al. (2019a, b); Pillepich et al. (2019) and the ones we utilized for calibration (red dots, detailed in Table 1).
In Fig. 3, we plot contours of ${\cal F}_{t}$ in the $\hat{\rho}_{H}-\hat{r}_{H}$ plane and overlay isolated TNG-50 halos and those used to calibrate the $\alpha$ and $\gamma$ parameters to delineate the parameter region of interest. Interestingly, the contours align well with the median behavior of the simulated halos. Halos that scatter towards the upper-right can have core collapse times several orders of magnitude shorter than their dark matter-only counterparts. Conversely, towards the lower-left, a large population of halos remains nearly unaffected by baryons.
The broad range of ${\cal F}_{t}$ values covered by the TNG-50 halos demonstrates that SIDM can significantly amplify diversity in inner halo structures. Since the SIDM dependence does not manifest in ${\cal F}_{t}$, our discussion is generic; physical results should be interpreted by factoring in the effective SIDM cross sections.
@@ III.3 The gravothermal phase of an SIDM halo with baryons
where the $dV_{\rm max,Model}(\tau^{\prime})/d\tau^{\prime}$ and $dR_{\rm max,Model}(\tau^{\prime})/d\tau^{\prime}$ terms are analytic equations taken from the parametric model that describe how fast the dark matter $V_{\rm max}$ and $R_{\rm max}$ evolve along the gravothermal phase $\tau(t)=t/t_{c}$. Based on the obtained $V_{\rm max}(t)$ and $R_{\rm max}(t)$, one can identify the fictitious CDM progenitor that gives rise to these values $\tau(t)$. This fictitious CDM halo only moderately differs from the CDM halo at time $t$; hence, replacing it with the CDM halo would only decrease a small amount of accuracy. In Ref. Yang et al. (2024b), a comparison of results with and without this simplification is performed for dark matter-only halos. In Appendix B, we verify this point through quantitative examples with baryons. Applications in this work will take advantage of this observation and use the CDM halo parameters to replace the fictitious ones.
@@
In Ref. Yang et al. (2024a), the integral approach is validated using halos from a cosmological SIDM simulation. The consistent agreement across halos of various mass scales and differing effective SIDM cross sections provides robust support for employing the idea of the gravothermal state. Notably, most merger events do not compromise the performance of the model because the changes in the gravothermal phases are negligible during these events. Even if a halo is in the collapsing phase during a merger, the effectiveness of the parametric model does not seem to decrease significantly. Such violation effect is explored in Ref. Yang et al. (2024b) using an extreme SIDM cross section simulation.
@@ VI Effect of SIDM on the baryon profiles
The obtained results are illustrated in the middle and right panels of Fig. 7. The middle panel case corresponds to the region below the dashed red curve. In this region, $r_{c,\rm max}>r_{H}$ and hence the baryon content can be more sensitive to the SIDM core evolution. In the simulated case, even if the stellar mass is reduced by a factor of $2^{3}$, the amount of size growth is comparable (slightly larger) to the case with the original $r_{H}$ value. The right panel case, instead, represents a case above the dashed red curve, where $r_{c,\rm max}<r_{H}$. This region is featured by a strong baryon effect on SIDM halos that can be cross-checked with contours in Fig. 3. In the case that we simulate, the total baryon mass is about $3.4$ times the original case and the core collapse time is reduced to about 3 Gyr. Interestingly, we found the baryon size evolution to have bypassed the expansion phase and go directly into the collapsing phase. This result can be correlated to the behavior in the dark matter distribution. For example, the DM12+baryon case has $\rho_{H}/\rho_{s}=149$ and $r_{H}/r_{s}=0.09$, which is well above the dashed red curve. Its inner density evolution corresponds to the magenta curve in Fig. 4, where we see no signs of core formation. Similarly, the DM11+baryon M2 case also resides above the dashed red curve and demonstrates no core formation.
@@ VII Applications
As with the DM-only parametric model, the baryonic extension of this model can be effectively applied to CDM simulations or semi-analytic models to forecast SIDM outcomes. The dark matter-only version of the model has recently be implemented into the Semi-Analytical Sub-Halo Inference ModelIng (SASHIMI) program Ando et al. (2024). Here, we demonstrate such an application using the Cored-DZ model on a sample of isolated halos. This application simplifies the modeling by considering generic accretion histories that do not differentiate between late and early-type galaxies. Additionally, we disregard the SIDM impacts on baryons discussed in Section VI, due to the absence of a calibration and validation using SIDM simulation data. Despite this omission, the influence on the maximum circular velocity ($V_{\rm max}$) and its corresponding radius ($r_{\rm max}$) is minimal when considering combined dark matter and baryon contributions. Thus, our results focus solely on these parameters for clarity.

### v2

@@ I Introduction
Cold dark matter (CDM) aligns well with observations at scales larger than megaparsecs. However, discrepancies arise at smaller scales in the inner structures of dark matter (DM) halos, where the density profiles can be either cored or cuspy Flores and Primack (1994); Spergel and Steinhardt (2000); de Blok et al. (2001); Gentile et al. (2004); Strigari et al. (2006); Donato et al. (2009); Kuzio de Naray et al. (2010); Oh et al. (2011); Oman et al. (2015); Salucci (2019); Di Paolo et al. (2019); De Laurentis and Salucci (2022). The cored cases may exhibit fainter or more diffuse profiles than most models predicted, while the cuspy cases can display steeper inner-density slopes. To explain these variations, researchers have elaborated on both baryonic feedback models and self-interacting dark matter (SIDM) models. Baryonic effects diversify the inner halo structures because it alters the structure by leading to compact stellar distributions through dissipation while simultaneously dispersing these distributions through feedback. The effectiveness of these processes varies with scale and depends on complex mechanisms including star formation and supernova feedback Navarro et al. (1996); Gnedin and Zhao (2002); Read and Gilmore (2005); Governato et al. (2010); Di Cintio et al. (2014); Chan et al. (2015); Tollet et al. (2016); Lazar et al. (2020); Macciò et al. (2020); Santos-Santos et al. (2020a). Similarly, SIDM contributes to diversity in inner halo structures through elastic scatterings that facilitate energy transfer within halos. This results in core formation and core collapse stages in the thermodynamic evolution of self-gravitating halos, known as gravothermal evolution Tulin and Yu (2018); Adhikari et al. (2022); Spergel and Steinhardt (2000); Kochanek and White (2000); Kamada et al. (2017); Ren et al. (2019); Santos-Santos et al. (2020b); Spergel and Steinhardt (2000); Correa et al. (2022); Yang et al. (2023a); Balberg et al. (2002, 2002); Koda and Shapiro (2011); Nishikawa et al. (2020); Sameie et al. (2020); Zavala et al. (2019); Kaplinghat et al. (2019); Kahlhoefer et al. (2019); Correa (2021); Turner et al. (2021); Slone et al. (2023); Silverman et al. (2022); Correa et al. (2022); Yang et al. (2023a); Minor et al. (2021); Yang and Yu (2021); Gilman et al. (2021, 2023); Nadler et al. (2023); Pollack et al. (2015); Choquette et al. (2019); Feng et al. (2021, 2022); Meshveliani et al. (2023); Fischer et al. (2024); Dhanasingham et al. (2023); Zhang et al. (2024); Tran et al. (2024). In probing SIDM, the diversity from baryon physics naturally integrate through gravitational interactions, resulting in a nonlinear and complex combined evolution. Nonetheless, understanding the interplay between baryons and dark matter is vital for making realistic SIDM predictions, as we depend on visible baryonic matter to probe the elusive dark matter structures Vogelsberger et al. (2014); Robertson et al. (2019, 2021); Yang and Yu (2021); Rose et al. (2023); Mastromarino et al. (2023); Rahimi et al. (2023); Zhong et al. (2023); Creasey et al. (2017); Fry et al. (2015); Kong et al. (2024); Fischer et al. (2023).
@@
To date, there have been only a handful of hydrodynamical cosmological simulations that incorporate SIDM physics Robertson et al. (2018); Robles et al. (2017); Elbert et al. (2018); Despali et al. (2019); Robertson et al. (2019). These simulations have revealed a complex interplay dependent on mass, galaxy type, and accretion history. Meanwhile, theoretical studies have used conducting fluid models, semi-analytic models, and idealized N-body simulations to model the combined effects of a baryonic potential and gravothermal evolution Balberg et al. (2002); Koda and Shapiro (2011); Essig et al. (2019); Nishikawa et al. (2020); Feng et al. (2021); Yang et al. (2023b); Jiang et al. (2023); Yang et al. (2023c); Zhong et al. (2023); Gad-Nasr et al. (2024). It has been illuminated that galaxy formation can adiabatically contract dark matter, increase the DM scattering rate, and thus accelerate the gravothermal evolution. Some empirical relations have also been proposed to enhance our understanding of these complex processes Elbert et al. (2018); Sameie et al. (2018); Feng et al. (2021); Yang et al. (2023c); Zhong et al. (2023).
@@
In Fig. 1, we highlight the diverse baryonic structures across a wide range of samples depicted on the plane of scale density ratio $\rho_{H}/\rho_{s}$ versus scale radius ratio $r_{H}/r_{s}$. Here, the halos and their stellar components are modeled using NFW and Hernquist Hernquist (1990) profiles, characterized by scale parameters $\rho_{s}$, $r_{s}$ for halos and $\rho_{H}$, $r_{H}$ for stellar components. The color codes on the plot correspond to the stellar-to-halo mass ratio, represented on a logarithmic scale, revealing a trend where higher $M_{s}/M_{h}$ halos move up along the $\ln(y)\propto\ln(x)$ direction, exhibiting larger $\rho_{H}/\rho_{s}$ and $r_{H}/r_{s}$ ratios. Red dots indicate benchmark halos from Table 1, utilized for validating the model. These halos cover a broad spectrum in the $\rho_{H}/\rho_{s}$ and $r_{H}/r_{s}$ plane, demonstrating the model’s broad applicability.
@@
Table 1: Properties of N-body simulated benchmarks: The first four simulations, introduced in Ref.Zhong et al. (2023), correspond to the initial block of cases in Table 1 therein. The final two benchmarks, DM13+baryon2 and DM13 extreme are analogous to J1636+4707 and JWST-ER1, as described in Ref. Kong et al. (2024). Additionally, the DM12+baryon, DM13 compact, and DM13+baryon benchmarks represent further simulations conducted to broaden the range of benchmark scenarios. The core collapse times are computed considering $\sigma/m=10~{}\rm cm^{2}/g$ for the DM11 cases in the first block and $\sigma/m=1~{}\rm cm^{2}/g$ for the DM12 and DM13 cases, to be consistent with the curves in Fig. 4. We use $t_{c,b}$ and $t_{c,0}$ to represent the core collapse times in the DM+baryon and DM-only scenarios. The halo masses are calculated as the mass enclosed within radii where the density is 200 times the critical density ($R_{\rm 200}$), using the scale density ($\rho_{s}$) and scale radius ($r_{s}$) of the NFW profile. The stellar masses are total masses computed using the scale density ($\rho_{H}$) and scale radius ($r_{H}$) of the Hernquist profile.
@@ III.2 A universal form factor for normalizing the evolution time with baryons
Figure 2: The unbroken universality in the evolution timeline with baryons, depicted through the evolution of normalized density profiles. For clarity, three benchmark cases—DM11 (blue), DM11+baryon (red), and DM13 compact (gray)—from Table 1 are illustrated to demonstrate the predictions of the parametric model. The $\rho_{s,0}$ and $r_{s,0}$ correspond to the initial NFW scale density and radius. The core collapse time $t_{c}$ is introduced in Eqs. (5) and (III.2) for the dark matter-only case and dark matter plus baryon case, respectively. All the cases evolve into a fixed configuration as $t/t_{c}$ approaching $1.1$. For more details, refer to Section V and Fig. 4.
Before delving into the specifics, we illustrate in Fig. 2 the aspects just discussed through the normalized evolution of density profiles with and without the presence of baryons. All three axes are normalized to depict universality and its breaking clearly. From this figure, it is evident that all scenarios follow a unified evolution timeline, represented as $t/t_{c}$ and ranging from 00 to $1.1$. We define each point along this timeline as a “gravothermal phase.” The density profiles, conversely, vary according to the amount of baryonic matter and show different deviations from the DM-only scenario. Subsequently, we will discuss methods to normalize the evolution time and adjust the DM density profile for a parametric model.
@@
Figure 3: Contours of ${\cal F}_{t}\equiv t_{c,b}/t_{c,0}$ on the $\rho_{H}/\rho_{s}$ ($\hat{\rho}_{H}$) versus $r_{H}/r_{s}$ ($\hat{r}_{H}$) plane (main panel), depicted together with the isolated halos in the TNG-50-1 simulation (black dots) Nelson et al. (2019a, b); Pillepich et al. (2019) and the ones we utilized for calibration (red dots, detailed in Table 1). The inset panel presents the median of a grid of ${\cal F}_{t}\equiv t{c,b}/t{c,0}$ over the $\hat{\rho}_{H}-\hat{r}_{H}$ plane, projected onto $M_{s}/M_{h}$ (red band), with the best fit shown in blue. Analogous results from the TNG-50-1 data are represented by black dots.
In Fig. 3 (main panel), we plot contours of ${\cal F}_{t}$ in the $\hat{\rho}_{H}-\hat{r}_{H}$ plane and overlay isolated TNG-50 halos and those used to calibrate the $\alpha$ and $\gamma$ parameters to delineate the parameter region of interest. Interestingly, the contours align well with the median behavior of the simulated halos. Halos that scatter towards the upper-right can have core collapse times several orders of magnitude shorter than their dark matter-only counterparts. Conversely, towards the lower-left, a large population of halos remains nearly unaffected by baryons. The broad range of ${\cal F}_{t}$ values covered by the TNG-50 halos demonstrates that SIDM can significantly amplify diversity in inner halo structures. Since the SIDM dependence does not manifest in ${\cal F}_{t}$, our discussion is generic; physical results should be interpreted by factoring in the effective SIDM cross sections.
In Fig. 3 (inset panel), we further display ${\cal F}_{t}$ as a function of $M_{s}/M_{h}$, calculated based on the properties of isolated halos in the TNG-50-1 simulation, as shown in Fig. 1. A clear correlation is observed between the two variables, with scatter increasing at higher $M_{s}/M_{h}$ values. To facilitate a rough estimate of ${\cal F}_{t}$ as a function of $M_{s}/M_{h}$, we derive a median relation from a fit, yielding the following expression:
${\cal F}_{t}(M_{s}/M_{h})\approx\frac{e^{-7.9(M_{s}/M_{h})}}{1+2300(M_{s}/M_{h})^{1.3}},$ (11)
which is applicable for $M_{s}/M_{h}\lesssim 0.08$.
@@ III.3 The gravothermal phase of an SIDM halo with baryons
where the $dV_{\rm max,Model}(\tau^{\prime})/d\tau^{\prime}$ and $dR_{\rm max,Model}(\tau^{\prime})/d\tau^{\prime}$ terms are analytic equations taken from the parametric model that describe how fast the dark matter $V_{\rm max}$ and $R_{\rm max}$ evolve along the gravothermal phase $\tau(t)=t/t_{c}$. Based on the obtained $V_{\rm max}(t)$ and $R_{\rm max}(t)$, one can identify the fictitious CDM progenitor that gives rise to these values $\tau(t)$. This fictitious CDM halo only moderately differs from the CDM halo at time $t$; hence, replacing it with the CDM halo would only decrease a small amount of accuracy. In Ref. Yang:2024uqb, a comparison of results with and without this simplification is performed for dark matter-only halos. In Appendix B, we verify this point through quantitative examples with baryons. Applications in this work will take advantage of this observation and use the CDM halo parameters to replace the fictitious ones.
@@
In Ref. Yang et al. (2024a), the integral approach is validated using halos from a cosmological SIDM simulation. The consistent agreement across halos of various mass scales and differing effective SIDM cross sections provides robust support for employing the idea of the gravothermal state. Notably, most merger events do not compromise the performance of the model because the changes in the gravothermal phases are negligible during these events. Even if a halo is in the collapsing phase during a merger, the effectiveness of the parametric model does not seem to decrease significantly. Such violation effect is explored in Ref. Yang:2024uqb using an extreme SIDM cross section simulation.
@@ VI Effect of SIDM on the baryon profiles
The obtained results are illustrated in the middle and right panels of Fig. 7. The middle panel case corresponds to the region below the dashed red curve. In this region, $r_{c,\rm max}>r_{H}$ and hence the baryon content can be more sensitive to the SIDM core evolution. In the simulated case, even if the stellar mass is reduced by a factor of $2^{3}$, the amount of size growth is comparable (slightly larger) to the case with the original $r_{H}$ value. The right panel case, instead, represents a case above the dashed red curve, where $r_{c,\rm max}<r_{H}$. This region is featured by a strong baryon effect on SIDM halos that can be cross-checked with contours in Fig. 3. In the case that we simulate, the total baryon mass is about $3.4$ times the original case and the core collapse time is reduced to about 3 Gyr. Interestingly, we found the baryon size evolution to have bypassed the expansion phase and go directly into the collapsing phase. This result can be correlated to the behavior in the dark matter distribution. For example, the DM12+baryon case has $\rho_{H}/\rho_{s}=149$ and $r_{H}/r_{s}=0.09$, which is well above the dashed red curve. Its inner density evolution corresponds to the magenta curve in Fig. 4, where we see no signs of core formation. Similarly, the DM11+baryon M2 case also resides above the dashed red curve and demonstrates no core formation. Our results align with a finding from Ref. Correa et al. (2024), which shows that some Milky Way mass galaxies, analogous to our DM12+baryon sample, tend to be more compact in SIDM.
@@ VII Applications
As with the DM-only parametric model, the baryonic extension of this model can be effectively applied to CDM simulations or semi-analytic models to forecast SIDM outcomes. The dark matter-only version of the model has recently been implemented into the Semi-Analytical Sub-Halo Inference ModelIng (SASHIMI) program Ando et al. (2024). Here, we demonstrate such an application using the Cored-DZ model on a sample of isolated halos. This application simplifies the modeling by considering generic accretion histories that do not differentiate between late and early-type galaxies. Additionally, we disregard the SIDM impacts on baryons discussed in Section VI, due to the absence of a calibration and validation using SIDM simulation data. Despite this omission, the influence on the maximum circular velocity ($V_{\rm max}$) and its corresponding radius ($r_{\rm max}$) is minimal when considering combined dark matter and baryon contributions. Thus, our results focus solely on these parameters for clarity.